Brief Exercise 11-4 Novak Company purchased machinery on January 1, 2017, for $97,600. The machinery is estimated to have a sNovak Company purchased machinery on January 1, 2017, for $97,600. The machinery is estimated to have a salvage value of $9,760 after a useful life of 8 years. Compute 2017 depreciation expense using the double-declining-balance method. Depreciation expense $ LINK TO TEXT Compute 2017 depreciation expense using the double-declining-balance method, assuming the machinery was purchased on October 1, 2017. (Round answer to 0 decimal places, e.g. 5,125.) Depreciation expense $ Click if you would like to Show Work for this question: Open Show Work
Answer #1

1) Double decline rate = 100/8*2 = 25%

2017 Depreciation expense = 97600*25% = 24400

2) 2017 Depreciation expense = 24400*3/12 = 6100